OpenStudy (anonymous):

The x-intercept of the line whose equation is 2x - 3y = 6 is i got -2?

OpenStudy (anonymous):

it should be 3 ...x intercept means y=0

OpenStudy (anonymous):

thats not right though

OpenStudy (amistre64):

define "not right"

OpenStudy (anonymous):

not the correct answer

OpenStudy (anonymous):

3 i right 2x - 0 = 6 x = 3

OpenStudy (anonymous):

it should be right ...it's x intercept O_o

OpenStudy (amistre64):

2x - 3y = 6 ; x int is y=0 2x - 3(0) = 6 2x = 6 x = 6/2 x = 3 ; now define "not the correct answer" please

OpenStudy (anonymous):

-2 is y intercept

OpenStudy (anonymous):

Maybe she's looking for a paired response (-2,0)

OpenStudy (anonymous):

Err (3,0)